# Query Rewriting via LLMs

### v3

@@ 5.1 MCTS Search Process
The pseudocode of the search procedure is shown in Algorithm 1. It consists of four stages – Selection, Expansion, Simulation and Back Propagation – that are repeated across $iter_{max}$ iterations:
1. Selection: The first stage is responsible for identifying the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure). Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1). Actions are picked using a UCB that balances exploration and exploitation. The goal is to pick those actions that have either (1) a higher potential to produce correct and faster rewrites (exploitation); or (2) been selected fewer times in the past (exploration).
@@
Here, $n^{\prime}$ is the node reached from $n$ by taking action $a$, and the first component $V(n^{\prime})$ represents the exploitation potential of $n^{\prime}$ to produce correct and faster queries (this notion is formalized below in Stage 3). The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often. Here, $P_{LLM}$ represents the next token probability, and $visits[n]$ is the number of times $n$ has been visited during the search process. $\beta$ is a function that controls the balance between exploration and exploitation. It depends on two hyperparameters $c_{base}$ and $c$ – a higher value of $c_{base}$ makes the algorithm favor exploitation , whereas a higher value of c increases the incentive to explore. $\beta$ is defined as:
@@
2. Expansion: The second stage is used to expand the unprocessed node $n_{cur}$ chosen by the Selection stage. Specifically, it retrieves from the LLM the top $k$ probable next tokens from $n_{cur}$’s state (Line 12), and expands the decision tree by adding $k$ new child nodes corresponding to these tokens. To make the expansion tractable, multiple child nodes are added only if the probability of the highest token falls below a threshold $\theta$ (Line 13). In other words, when the highest token probability is below $\theta$, it means that the LLM itself is unsure of what the next token should be and therefore it is worth exploring additional options. On the other hand, if the highest token probability is greater than $\theta$, then the tokens are generated in a greedy fashion from the current node until a point where the LLM is again unsure of the next token, or it reaches a terminal node (i.e., completes a query rewrite).
3. Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$. This node is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21). Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23). For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides w.r.t the original query. However, if invalid (i.e. syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value. After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
@@ 6.1 LITHE Parameter Settings
Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2).
@@ 6.2 Query Equivalence Testing
We use a multi-stage approach, described below, to help the DBA test semantic equivalence between the original query and a recommended rewrite. Note that, as mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems[14].
@@
2. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is NP-complete [7], a variety of logic-based tools (e.g. Cosette[12], SQL-Solver [17], VeriEQL [21], QED [40]) are available for proving equivalence over restricted classes of queries. Once the least-cost rewrite is identified, LITHE uses the recently proposed QED [40] and SQLSolver [17] since these two approaches were found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome. Note that this rewrite has already passed the sampling-based tests described above.
@@ 6.3 Regression Identification
We use heuristics to predict whether the actual execution time of a given promising rewrite may turn out to be slower than the original query. The first heuristic is based on the SEER robust plan identification algorithm [19], which is predicated on the database engine providing a “force-plan” feature (e.g. [1]). Let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively. We construct parametrized versions of these queries, where the constants in the filter predicates are replaced by variables. Then, by assigning appropriate values to these parameters, we construct queries that are located at the corners of the selectivity space. The plans $P^{U}$ and $P^{T}$ are forced at each of these corner locations, and if the rewrite’s cost is lower than the original at all of them, the rewrite is deemed to be robust.
At this time, plan forcing is not available in all database engines. For such limited environments, we fall back to an alternative heuristic – we compare the query runtimes obtained on the series of executions carried out on sampled databases during semantic equivalence testing. If the rewrite’s runtimes are lower in all of these intermediate executions, then it is estimated to be robust.
@@ 7 Experimental Evaluation
LITHE has also been evaluated on other benchmarks, including DSB [15], ARCHER [47], JOB [23] and StackOverflow [30]. The overall performance characteristics were found to be similar to those presented here – due to space limitations, we defer their details to Section A.
Testbed. The majority of our experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA. Variations on this platform are considered in Sections 7.4 and  7.5.
Metrics. For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with $>$ 1.5 speedup) could be constructed. Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e. CPRs arising from either LITHE or SOTA).
@@ 7.1.1 Estimated Cost
LITHE produces a rewrite with a positive cost speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold. Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5. On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1. All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27. Of these 27, we were able to formally verify 11 using the logic-based tools, whereas the remaining 16 passed our statistical tests. Furthermore, we also manually verified the correctness of these rewritten queries.
@@
We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE. This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
@@
Second, in most queries, LITHE’s cost speedup either exceeds or matches SOTA. In fact, for several queries (e.g. Q45, Q25, Q4), LITHE produces a highly beneficial CPR but SOTA returns the original query. Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query. And in Q88 and Q95, SOTA performs only marginally better.
At this stage, one might expect that adding more rules to LITHE could bring it on par with SOTA for queries like Q57. However, we deliberately include only broad-brush rules in LITHE to ensure generalizability and efficiency. As the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) in terms of runtime.
@@ 7.1.2 Execution Time
Thus far, we had considered optimizer-estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA on a cold-cache environment. We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a large factor of 700! Second, in all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well. On the one hand, there is Q45 where the projected speedup of 40X increases to 700X at runtime. When run in a warm-cache environment, this speedup increases to a huge $10^{4}$ (see Section A). On the other hand, the $10^{5}$ speedup for Q41 decreases to a still reasonable 200X. But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100. However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites. Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA provides a TSGM of 4.9.
@@ 7.1.3 Reasoning
As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries. For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700X speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc. Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
@@ 7.2.1 Components of LITHE
We observe that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) contribute to around two-thirds of the CPRs (18/26). When the statistics-infused Rules (R5, R6) are added, this number jumps to 25. Finally, MCTS produces an additional CPR and also improves the cost speedup of a pre-existing CPR.
@@
Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance. The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads. When this experiment was conducted, the CSGM dropped precipitously to a paltry 5, a far cry from the 11.5 obtained with the database-sensitive rules. These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
@@ 7.3.1 Transformation Time
The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes. However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in just 3 minutes. Further, many applications tend to use a set of canned queries which are run thousands of times. Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
@@
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module. We explore techniques to improve their efficiency in Section B.

### v4

@@ 5.1 MCTS Search Process
The search procedure is codified in Algorithm 1, and comprises four stages – Selection, Expansion, Simulation and Back Propagation – described below, which are repeated for $iter_{max}$ iterations.
1. Selection: The first stage identifies the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure). Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1). Actions are picked using a UCB that balances exploration and exploitation. The goal is to pick those actions that have either (1) a higher potential to produce correct and faster rewrites (exploitation); or (2) been selected fewer times in the past (exploration).
@@
Here, $n^{\prime}$ is the node reached from $n$ by taking action $a$, and the first component $V(n^{\prime})$ represents the exploitation potential of $n^{\prime}$ to produce correct and faster queries (this notion is formalized below in Stage 3). The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often. Here, $P_{LLM}$ represents the next token probability, which can be queried from the LLM by enabling the logprobs option in the API, and $visits[n]$ is the number of times $n$ has been visited during the search process. $\beta$ is a function that controls the balance between exploration and exploitation. It depends on two hyperparameters $c_{base}$ and $c$ – a higher value of $c_{base}$ makes the algorithm favor exploitation , whereas a higher value of c increases the incentive to explore. $\beta$ is defined as:
@@
Figure 7: Sample decision tree traversal (Algorithm 1)
For example, consider the state of the tree in Figure 7 when Nodes 5 and 15 are the current unexpanded nodes. At this juncture, the selection procedure will use the UCB values of these two nodes to choose which node to expand next.
2. Expansion: The second stage is used to expand the unprocessed node $n_{cur}$ chosen by the Selection stage. Specifically, it retrieves from the LLM the top $k$ probable next tokens from $n_{cur}$’s state (Line 12), and expands the decision tree by adding $k$ new child nodes corresponding to these tokens. To make the expansion tractable, multiple child nodes are added only if the probability of the highest token falls below a threshold $\theta$ (Line 13). Otherwise, the tokens are generated in a greedy fashion from the current node until a point where the LLM is again unsure of the next token, or it reaches a terminal node (i.e. completes a query rewrite).
In our implementation, $k$ and $\theta$ are set to 2 and 0.7, respectively. Since branching occurs only when this relatively high $\theta$ threshold is breached, most nodes have only one child, resulting in long linear sequences. This design, combined with at most two nodes being processed during a branch, significantly reduces the effective fan-out of the tree. Moreover, it ensures that MCTS focuses computational resources on the brittle segments of the rewrite, and does not waste them on robust modifications. For example, when the root node in Figure 7 is first processed, then Nodes 1, 2, 3 and 4 are expanded and created sequentially since the token probabilities are higher than $\theta$ (Lines 24–26 of the while loop). Only when Node 4 is reached, two new nodes (5 and 15) are created as its children.
3. Simulation: The node $n_{cur}$ is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21). The rewritten query obtained from this greedy expansion is then used to compute the potential $V(n_{cur})$ (Line 22) as follows: for a valid rewrite, $V(n_{cur})$ represents the speedup it provides w.r.t. the original query. However, if invalid (i.e. syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value. After every simulation, the rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$ (Line 23).
For instance, to compute $V(n_{15})$ in Figure 7, the blue path in the tree is greedily expanded to identify its rewrite potential. That is, the partial rewrite “select * from B left outer join A;” is used to evaluate $V(n_{15})$.
@@ 5.3 Performance
Discussion. Our approach implicitly assumes that LLMs do possess some degree of knowledge about query optimization – for example, when presented with a poorly written query versus an optimized one, the LLM is usually capable of identifying why the optimized version would execute better. This expertise arises from the extensive training corpus of LLMs, which includes a large repository of books, papers, and articles on query optimization and execution.
However, this core knowledge is not fully sufficient to provide good rewrites, and there is a clear need for augmentation with explicit query-specific guidance. This is particularly so when there are multiple plausible optimization paths, or when the LLM may be subtly misled – this is the reason we leverage MCTS to systematically explore alternatives based on token probabilities. Further, there is explicit evaluation via the database optimizer of the expected improvement in quantitative performance. This separation ensures that our method leverages the LLM’s generative capabilities to posit rewrites while relying on an external ground-truth signal (i.e. the database optimizer) to confirm rewrite quality.
@@ 6.1 LITHE Parameter Settings
Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2) – if this threshold is crossed, we drop the rewrite.
@@ 6.2 Query Equivalence Testing
We use a multi-stage approach, described below, to help the DBA test semantic equivalence between the original query and a recommended rewrite. As mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems [17].
@@
2. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is NP-complete [10], a variety of logic-based tools (e.g. Cosette[15], SQL-Solver [20], VeriEQL [26], QED [49]) are available for proving equivalence over restricted classes of queries. Once the least-cost rewrite is identified, LITHE uses QED [49] and SQLSolver [20] since these two approaches covered a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome. Note that this rewrite has already passed the sampling-based tests described above.
@@ 6.3 Regression Identification
We use heuristics to predict whether the actual execution time of a given promising rewrite may turn out to be slower than the original query. The first heuristic is based on the SEER robust plan identification algorithm [22], which is predicated on the database engine providing a “force-plan” feature (e.g. [1]). This approach is based on the notion of replacement safety, wherein the replacement plan is always either better or within a small sub-optimality factor of the original optimizer-recommended choice. It was theoretically shown that a safe plan replacement at the perimeter of the selectivity space is sufficient to infer safe replacement in the interior of the space as well. The proof used a generic characterization of plan cost functions and the first and second derivatives of these functions at the perimeter. As a matter of efficiency, it was empirically found that instead of the entire perimeter, it was usually sufficient to consider safety just at the corners.
We use the above result to design a regression test as follows: let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively. We construct parametrized versions of these queries, where the constants in the filter predicates are replaced by variables. Then, by assigning appropriate values to these parameters, we construct queries that are located at the corners of the selectivity space. The plans $P^{U}$ and $P^{T}$ are forced at each of these corner locations, and if the rewrite’s cost is lower than the original at all of them, the rewrite is deemed to be robust.
At this time, plan forcing is not available in all database engines. For such limited environments, we fall back to an alternative heuristic – we compare the query runtimes obtained on the series of executions carried out on sampled databases during semantic equivalence testing. If the rewrite’s runtimes are lower in all of these intermediate executions, then it is estimated to be robust. We found this extrapolation heuristic to be effective in our experiments due to two reasons: First, the sampled databases, while relatively small compared to the original, are still large in absolute terms (1 GB or more). Second, since most rewrites are related to redundancy removals, a relative performance improvement obtained due to such a rule holds true irrespective of the data sizes.
@@ 7 Experimental Evaluation
LITHE has also been evaluated on other benchmarks, including DSB [18], ARCHER [56], JOB [29] and StackOverflow [36]. The overall performance characteristics were found to be similar to those presented here – see Section A for details.
Testbed. The experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660, 32 GB RAM, and 12 TB HDD, running Ubuntu 22.04 LTS. A majority of the experiments used PostgreSQL v16 database engine and GPT-4o LLM for both LITHE and SOTA. Commercial engines and other LLMs are considered in Sections 7.4 and  7.5.
Metrics. For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with ¿ 1.5 speedup) could be constructed. Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e. CPRs arising from either LITHE or SOTA). Cost speedups are computed relative to the native optimizer, and runtime speedups are measured as the ratio of original query runtime to rewritten query runtime in a cold-cache environment.
@@ 7.1.1 Estimated Cost
LITHE produces a rewrite with a positive cost speedup (¿ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold. Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5. On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1. All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27. Of these 27, we were able to formally verify 11 using the logic-based tools, whereas the remaining 16 passed our statistical tests. Furthermore, we also manually verified the correctness of these rewritten queries.
@@
We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups over the native engine – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE. This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter efficiently checks result existence in an inner subquery since it removes the costly aggregation function.
@@
Second, in as many as 15 queries, LITHE’s cost speedup substantively exceeds SOTA, while in the remainder it largely matches SOTA. In fact, for several queries (e.g. Q45, Q25, Q4) LITHE produces a highly beneficial CPR but SOTA returns the original query. The only case where SOTA is appreciably better is Q57 where it projects a large speedup but LITHE settles for the original query. And in Q88 and Q95, SOTA performs marginally better.
At this stage, one might expect that adding more rules to LITHE could bring it on par with SOTA for queries like Q57. However, we deliberately include only broad-brush rules in LITHE to ensure generalizability and efficiency. Moreover, as the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) at runtime!
We also analyzed queries where LITHE was unable to provide CPRs to determine whether there were common patterns that led to unsuccessful outcomes. Our audit indicated that the primary reasons were: (a) Structural Simplicity – for instance, flat SPJ queries without nesting that are already conducive to current optimizers (e.g. Q3, Q15, Q52), or (b) Structural Tightness – for instance, queries without repeated computations where the scope for improvement is inherently limited (e.g. Q7, Q19, Q22). Further, some failures may be due to the specific LLM being used, and not the queries themselves (see Section 7.5.2). Therefore, one could consider implementing an ensemble of diverse LLMs to gain enhanced beneficial coverage.
@@ 7.1.2 Execution Time
Thus far, we had considered optimizer-estimated costs. We now move to wall-clock runtimes for query executions – Figure 9 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA. We again observe that there are indeed several queries where substantial time benefits are achieved by the rewrites over the native engine, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a huge factor of 700! Further, for 12 of the 26 queries, LITHE significantly outperforms SOTA – as a case in point, by 14 times in Q11. In the remainder, it matches SOTA, including as mentioned above, queries where SOTA’s optimizer costs were better.
From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well. On the one hand, there is Q45 where the projected speedup of $40$ increases to $700$ at runtime, whereas on the other, the $10^{5}$ speedup for Q41 decreases to $200$. But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100. However, the good news is that with LITHE, although the runtime speedups did not always match the projections, regressions were not encountered among the CPR rewrites thanks to the sampling-based checks. Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA only delivers TSGM of 4.9.
@@ 7.1.3 Reasoning
As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries. For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700700x speedup as “the use of CTEs”, “Reduction in data volume early on” due to pre-filtering based on join predicates, etc. Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
@@ 7.2.1 Components of LITHE
We observe that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) contribute to around two-thirds of the CPRs (18/26). When the statistics-infused Rules (R5, R6) are added, this number jumps to 25. Finally, MCTS in isolation (with Prompt 1 as the seed prompt) produces only 11 CPRs. However, by using the best prompt as seed, an additional CPR is obtained and the cost speedup of a pre-existing CPR is also improved.
@@
Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance. The motivation is that it would relieve us from using the database-sensitive rules R1–R6 which incur significant computational and financial overheads. When this experiment was conducted, the CSGM dropped precipitously to a paltry 5, a far cry from the 11.5 obtained with the database-sensitive rules. These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
@@ 7.3.1 Transformation Time
We first look at the end-to-end time required to perform a rewrite (i.e. to run the entire pipeline shown in Figure 3). Table 6 shows the average processing time per CPR query, where we see that the LITHE rewrite process takes a few minutes. However, note that this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in under 3 minutes. Further, many applications tend to use a set of canned queries which are run thousands of times. Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
@@
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down showed the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module. We explore techniques to improve their efficiency in Section B.